OpenStudy (anonymous):

An investor invested a total of $2,000 in two mutual funds. One earned a 6% profit while the other earned 3% profit. If the investors total profit was $99 how much was invested in each mutual fund?

OpenStudy (anonymous):

@amistre64 @uri @zzr0ck3r

OpenStudy (anonymous):

@CarlosGP im lost

OpenStudy (anonymous):

ok i understand that @CarlosGP but, how much was invested into each fund?

OpenStudy (anonymous):

@CarlosGP i can solve it

OpenStudy (anonymous):

@CarlosGP nope,@heybhai solve it

OpenStudy (anonymous):

x+y=2000......eq1 0.06x+0.03y=99 =>6x+3y=9900 =>2x+y=3300.....eq2 subtract eq1 from eq2 we get 2x+y-(x+y)=3300-2000=>x=1300 put the value of x in eq1 we get y=2000-1300=>y=700 so $1300 were invested in 6% profit making MF and rest $70o in other one

OpenStudy (anonymous):

@CarlosGP u blocked me roflllllllllllllllllllllllll this aint facebook goofy u r funny and childish for that im looking 4 help idk what u r on????

OpenStudy (anonymous):

exception?????? y r u even still on this post get lost r u looking for some attention im asking 4 help if u dont want 2 help get lost i dont go back and forth over the www goofy